OpenStudy (anonymous):

can someone please explain how you work this.. 7w+2=3w+94

OpenStudy (anonymous):

first isolate w 7w-3w=94-2 4w=92 then divide both sides by 4 so u get only one w w=23

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

your welcome ;)